An airplane flew for 3 hours with a tail wind of 18 kilometers per hour. The return flight against the wind took 4 hours. Find the rate of the plane in still air.
Answer by mananth(16946)

plane speed =x mph
wind speed 18 mph
time with tail wind 3 hours
time against wind 4 hours
Distance = same
8(x+18) =4(x-18)
8x +144 =4x-72
4x-8x=144-72
x =216 mph plane speed